# iSeg: Interactive 3D Segmentation via Interactive Attention

### v1

@@ 4.1. Generality of iSeg
Table 1. Perceptual user study. We compare the 3D segmentation effectiveness for different clicked-based interactive methods on a scale of 1 to 5, corresponding to completely ineffective and completely effective segmentation. Our method is considered much more effective than the competitors.
4.2. Fidelity of iSeg
Our method’s training is supervised by segmentation masks generated from SAM (Kirillov et al., 2023) for 2D renderings of the shape. As we show in Figs. 3 and 9, SAM’s masks differ substantially between views. In contrast, iSeg manages to fuse the noisy training examples into a coherent 3D segmentation model that corresponds to the clicked vertices. Examples are presented in Figs. 2 and 6.
In Fig. 6, we further show the results of iSeg for a couple of clicks. We incorporate either a second positive click that extends the segmented part or a second negative click that retracts the region. For example, with the first click, the fine-grained bulb area of the lamp is segmented. Then, the second positive click enables including the flat surface surrounding the bulb. The negative click, on the other hand, offers the control to reduce and refine the segmentation region. As shown in Fig. 6, the first click on the hammer’s head segments the entire head. The front part can be easily and intuitively removed by a second negative click on it.
Perceptual user study. As far as we can ascertain, there is no dataset with ground-truth annotation for click-based interactive segmentation of 3D shapes. Thus, for quantitative evaluation of our method, we opt to perform a perceptual user study. We used 20 meshes from different categories, such as humanoids, animals, and man-made objects, and included 40 participants in our study.
@@ 4.2. Fidelity of iSeg
For comparison, we considered two alternatives, InterObject3D (Kontogianni et al., 2023) and a baseline we constructed based on SAM’s 2D segmentations. InterObject3D is a recent work on interactive segmentation of 3D objects. For the SAM baseline, we rendered the shape and projected the clicked point to 2D from 100 random views, computed SAM’s mask, and re-project the result back to 3D for each visible vertex. Then, we averaged the predictions according to the number of times each vertex was seen. Other interactive segmentation techniques use an implicit 3D representation (Chen et al., 2023b) or perform a different task (object detection) (Zhang et al., 2023), and thus, are not directly comparable to our method.
Fig. 15 presents several example questions from our study, demonstrating the comparison between the alternative segmentation techniques, and Tab. 1 summarizes the effectiveness score averaged over all the meshes and participants in the study. Fig. 13 further compares our method with the SAM baseline. As seen in the figures and reflected by the table, the participants rate the effectiveness of iSeg much higher than the other methods, indicating its fidelity to the clicked point by the user.
Applications. iSeg computes contiguous and localized shape partitions. These segmentations enable geometric local editing of the mesh, as demonstrated in Fig. 11. Further details are provided in the supplementary material.
4.3. Properties of iSeg
Stability. A desired property of a click-based segmentation model is to be stable. That is, the results should be similar for nearby clicks that relate to the same shape part. This property implies that the model is robust to the exact click location and alleviates the need for precision from the user. Fig. 7 demonstrates the stability of our method. As the figure shows, iSeg consistently selects the handle part of the hammer for different clicks along the handle.
We evaluate the stability property quantitatively by comparing the segmentation prediction for neighboring vertices. Given a clicked point, we compute the Intersection over Union (IoU) for its resulting mask and the mask for clicking each of its one-ring neighbors. We use 7 meshes with 100 random clicks from each mesh. For comparison, we utilize the SAM baseline, as described in Sec. 4.2. The average IoU for iSeg is 0.8 compared to 0.7 for the SAM baseline. SAM’s masks originate from 2D views of the shape. These predictions do not segment occluded regions of the shape and are less smooth in 3D. In contrast, our method operates directly in 3D and exhibits better stability than the SAM baseline.
[…]
@@ 4.3. Properties of iSeg
3D consistency. Different from SAM (Kirillov et al., 2023), our method operates natively in 3D. iSeg predicts the segmentation probability at the vertex level, or in other words, our result lives on the 3D surface of the mesh. Thus, it is 3D-consistent by construction. In Fig. 9, we compare our segmentation projected to 2D with SAM’s result for the projected shape views. Since SAM operates on each view independently, it yields inconsistent predictions. By contrast, in our case, any 2D projection originates from the same 3D segmentation and thus, our method is consistent across all views.
Generic feature information. Our mesh feature field is distilled directly from SAM’s encoder and is independent of the user’s point clicks for segmentation. Thus, although iSeg is optimized per mesh, the semantic feature representation is shared across shapes. We demonstrate this property by cross-domain segmentation. In this experiment, we use the point encoded features from one shape and predict region probability with a different shape’s decoder. Fig. 8 shows examples. The transferable features enable the creation of cross-domain shape analogies, such as how the belly of a human corresponds to the “belly” of an airplane.
4.4. Generalization Capabilities
Unseen vertices. Our method exhibits strong generalization power. First, we emphasize that we train just on a small fraction of 3% of the mesh vertices. Still, iSeg is successfully applied to other vertices unseen during training and properly respects the clicked points, as shown in Figs. 2 and 6 and discussed in Sec. 4.2. We note that all the results shown in the paper are for test vertices.
[…]
Figure 10. Customized segmentations. iSeg is capable of creating customized segmentations specified by several input clicks.
Unseen views. Although iSeg was trained with 2D supervision only, its predictions are 3D in nature. Sec. 3.2 exemplifies this phenomenon. A click on the back side of the backrest segments the front side as well. Such supervision does not exist for our model’s training, since the clicked vertex cannot be seen from the front side. Similarly, two clicks at opposite sides of the backrest segment the entire part, although they cannot be seen together from any single 2D view. This result suggests that iSeg learned 3D-consistent semantic vertex information, enabling it to generalize beyond its 2D supervision.
Unseen number of clicks. For resource efficient training, we trained iSeg on up to two clicks: single click, second positive click, and second negative click. Nonetheless, our model offers customized segmentation with more than two clicks, as demonstrated in Fig. 10. We attribute this capability to the interactive attention mechanism. The interactive attention layer seemed to learn the representation of a positive and a negative click, and how to attend to each click for a meaningful multi-click segmentation.
@@ 5. Conclusion
In this work, we presented iSeg, a technique for interactively generating fine-grained tailored segmentations of 3D meshes. We opt to lift features from a powerful pre-trained 2D segmentation model onto a 3D mesh, which can be used to create customized user-specified segmentations. Our mesh feature field is general and can be used for additional tasks such as cross-segmentation across shapes of different categories (e.g. Fig. 8). Key to our method is an interactive attention mechanism that learns a unified representation for a varied number of positive or negative point clicks. Our 3D-consistent segmentation enables selecting points across occluding surfaces and segmenting meaningful regions directly in 3D (e.g. Sec. 3.2).
In the future, we are interested in exploring additional applications of mesh feature fields beyond segmentation. We have demonstrated they may be used for cross-domain segmentation, and there may be other exciting applications, such as key-point correspondence, texture transfer, and more.
@@
The following sections provide more information regarding our interactive segmentation method. Appendix A shows additional results and experiments we conducted with iSeg. Appendix B discusses our perceptual user study. In appendix C, we detail the results of an ablation test. Finally, in appendix D, we elaborate on our implementation details for the encoder and decoder networks and the comparison settings to previous work.
@@ Appendix A Additional Results
View generalization, cross-segmentation, and stability. Figs. 16, 17 and 20 present additional examples of the view generalization of iSeg, its cross-domain segmentation ability, and the stability of the selected region. In Fig. 16, we show that iSeg segments the whole body of the guitar, even though its back side is occluded from the click on the front. Moreover, the second click on the back side of the guitar’s bridge cannot be seen together in any 2D view with the first click. Nevertheless, our method selects the complete 3D regions of the bridge and body.
Fig. 17 shows that the semantic information of iSeg’s features is shared between different shapes, and enables cross-domain segmentation of similar parts. Interestingly, for the click on the front right leg of the camel, the front right leg of the chair was mostly selected.
In Fig. 20, we present the segmentation stability of iSeg. For a couple of clicks, the user can select the entire trailer box or only its side with various locations for the positive clicks or positive and negative clicks, respectively.
Granularity. Fig. 18 exemplifies an additional property of our method: segmentation granularity. When segmenting the leg for an intricate shape like the human body, the user can control whether to exclude the foot, the lower shin part, or the upper part of the calf muscle by simply clicking the negative point at different locations along the leg.
Separability. We analyze iSeg’s segmentation prediction by plotting the histogram of probability values for all the shape vertices for different clicks. Results are presented in Fig. 21. We observe that the histogram contains two populations: vertices with high segmentation probability, which correspond to the region for the clicked point, and the rest of the vertices with low probability. To further demonstrate the separability of the model’s predictions, we use the Otsu threshold and binarize the probabilities. As shown in the figure, the region before and after thresholding remains almost the same, indicating that the segmented part can be easily separated from the shape.
@@
Figure 17. Cross-domain segmentation. iSeg learned features generalize across domains. The feature vector of a clicked point results in the selection of a similar semantic region for the shape itself and a shape from a different domain.

### v2

@@ 4.1. Generality of iSeg
Table 1. Quantitative evaluation on PartNet. We compare the segmentation performance of different clicked-based interactive techniques on shapes from the PartNet dataset (Mo et al., 2019). IoU stands for Intersection over Union. iSeg’s scores are substantially higher than those of the alternatives.
Method InterObject3D SAM Baseline iSeg (ours)
Accuracy $\uparrow$ 0.54 0.76 0.95
IoU $\uparrow$ 0.38 0.51 0.90
4.2. Fidelity of iSeg
Our method’s training is supervised by segmentation masks generated from SAM (Kirillov et al., 2023) for 2D renderings of the shape. As we show in Fig. 3, SAM’s masks differ substantially between views. In contrast, iSeg manages to fuse the noisy training examples into a coherent 3D segmentation model that corresponds to the clicked vertices. Examples are presented in Figs. 2 and 6. In the supplementary, we further demonstrate the method’s 3D consistency.
In Fig. 6, we further show the results of iSeg for a couple of clicks. We incorporate either a second positive click that extends the segmented part or a second negative click that retracts the region. For example, with the first click, the fine-grained bulb area of the lamp is segmented. Then, the second positive click enables to include the flat surface surrounding the bulb. The negative click, on the other hand, offers the control to reduce and refine the segmentation region. As shown in Fig. 6, the first click on the hammer’s head segments the entire head. The front part can be easily and intuitively removed by a second negative click on it.
Quantitative evaluation
As far as we can ascertain, there are no annotated datasets for click-based interactive segmentation of 3D shapes. Thus, we adapted the part segmentation dataset PartNet (Mo et al., 2019) for our setting. The evaluation included 170 meshes sourced from all the categories in the dataset. For each mesh, we selected five test vertices from a part at random, where each vertex was regarded as a single click, and measured how well the part was segmented. We used two evaluation metrics: Accuracy (ACC) and Intersection over Union (IoU). Further details about the evaluation are provided in the supplemental material.
We considered two alternatives for comparison, InterObject3D (Kontogianni et al., 2023) and a baseline we constructed based on SAM’s 2D segmentations. InterObject3D is a recent work on interactive segmentation of 3D objects. In our experiments, we employed the publicly available pretrained model released by the authors.
For the SAM baseline, we rendered the shape and projected the clicked point to 2D from 100 random views, computed SAM’s mask, and re-projected the result back to 3D for each visible vertex. Then, we averaged the predictions according to the number of times each vertex was seen. In the supplementary material, we discuss additional baselines we devised using SAM. Other interactive segmentation techniques use an implicit 3D representation (Chen et al., 2023b) or perform a different task (object detection) (Zhang et al., 2023), and thus, are not directly comparable to our method.
Tab. 1 presents the ACC and IoU averaged over the test clicks and shapes, and Fig. 7 shows visual examples. InterObject3D does not select the part properly and yields a partial segmentation. SAM Baseline segments the region of the part where the click is visible. However, occluded regions that belong to the part are not marked. In contrast, iSeg adheres to the clicked point. It segments a coherent region in 3D, which is similar to the ground-truth part label, and achieves much higher ACC and IoU than the baselines.
Perceptual user study
iSeg is not limited to a particular shape type from a dataset or specific parts defined in the dataset. In such cases, ground-truth labels are unavailable. Thus, to evaluate the effectiveness of the flexible and diverse segmentations offered by our method, we opt to perform a perceptual user study. We used 20 meshes from different categories, such as humanoids, animals, and man-made objects, and included 40 participants in our study.
Table 2. Perceptual user study. We evaluate the 3D segmentation effectiveness on a scale of 1 to 5, corresponding to completely ineffective and completely effective segmentation. Our method is considered much more effective than the competitors.
@@ 4.2. Fidelity of iSeg
Tab. 2 summarizes the effectiveness score averaged over all the meshes and participants. Fig. 15 further compares our method with the SAM baseline. As seen in the figure and reflected by the table, the participants rated the effectiveness of iSeg much higher than the other methods, indicating its fidelity to the clicked point.
Applications
iSeg computes contiguous and localized shape partitions. These segmentations can be used to extract shape parts easily and enable applications such as full-shape segmentation and local geometric editing of the mesh. These results are demonstrated in Figs. 9, 10 and 13. Further details and discussion appear in the supplementary material.
4.3. Generic Feature Information
Figure 7. Segmentation comparison on PartNet. Each pair shows different views of the segmentation result for different interactive techniques. The other methods produce a partial segmentation of the part containing the click. In contrast, iSeg obtains an accurate result, which is similar to the ground-truth annotation of the part.
Our mesh feature field is distilled directly from SAM’s encoder and is independent of the user’s clicks for segmentation. Thus, although iSeg is optimized per mesh, the semantic feature representation is shared across shapes. We demonstrate this property by cross-domain segmentation. In this experiment, we use the point encoded features from one shape and predict region probability with a different shape’s decoder. Fig. 8 shows examples. The transferable features enable the creation of cross-domain shape analogies, such as how the belly of a human corresponds to the “belly” of an airplane.
4.4. Generalization Capabilities
Unseen mesh vertices
Our method exhibits strong generalization power. First, we emphasize that we train just on a small fraction of 3% of the mesh vertices. Still, iSeg is successfully applied to other mesh vertices unseen during training and properly respects the clicked points, as shown in Figs. 2 and 6 and discussed in Sec. 4.2. We note that all the results shown in the paper are for test vertices.
@@ 4.4. Generalization Capabilities
Unseen views
Although iSeg was trained with 2D supervision only, its predictions are 3D in nature. Fig. 5 exemplifies this phenomenon. A click on the back side of the backrest segments the front side as well. Such supervision does not exist for our model’s training, since the clicked vertex cannot be seen from the front side. Similarly, two clicks at opposite sides of the backrest segment the entire part, although they cannot be seen together from any single 2D view. This result suggests that iSeg learned 3D-consistent semantic vertex information, enabling it to generalize beyond its 2D supervision.
Unseen number of clicks
For resource efficient training, we trained iSeg on up to two clicks: single click, second positive click, and second negative click. Nonetheless, our model offers customized segmentation with more than two clicks, as demonstrated in Fig. 11. We attribute this capability to the interactive attention mechanism. The interactive attention layer seemed to learn the representation of a positive and a negative click, and how to attend to each click for a meaningful multi-click segmentation.
Limitations
Our method may not follow the symmetry of the mesh exactly, as exemplified in Fig. 12. For a click on the goat’s head, the segmented regions from the sides of the head differ somewhat from each other, since iSeg is not trained to segment those regions the same. In the supplementary, we discuss the potential limitation of processing 3D shapes using MLPs operating in the Euclidean space.
@@ 5. Conclusion
In this work, we presented iSeg, a technique for interactively generating fine-grained tailored segmentations of 3D meshes. We opt to lift features from a powerful pre-trained 2D segmentation model onto a 3D mesh, which can be used to create customized user-specified segmentations. Our mesh feature field is general and may be used for additional tasks such as cross-segmentation across shapes of different categories (e.g. Fig. 8). Key to our method is an interactive attention mechanism that learns a unified representation for a varied number of positive or negative point clicks. Our 3D-consistent segmentation enables selecting points across occluded surfaces and segmenting meaningful regions directly in 3D (e.g. Fig. 5).
In the future, we are interested in exploring additional applications of iSeg beyond segmentation. We have demonstrated that it can potentially be used for cross-domain segmentation, and there may be other exciting applications, such as key-point correspondence, texture transfer, and more.
Acknowledgements.
This research was supported by grant #2022363 from the United States - Israel Binational Science Foundation (BSF), grant #2304481 from the National Science Foundation (NSF), and gifts from Adobe, Snap, and Google. We thank the University of Chicago and the Toyota Technological Institute at Chicago (TTIC) for allocating computational resources for this work, and their technical staff for the support along the project. We also extend our gratitude to the members of the 3DL lab at the University of Chicago for their helpful feedback and excellent advice.
@@ Acknowledgements.
The following sections provide more information regarding our interactive segmentation method. Appendix A shows additional results and experiments we conducted with iSeg. Appendices B and C discuss our quantitative evaluation on the PartNet dataset and the perceptual user study, respectively. In appendix D, we detail the results of an ablation test. Finally, in appendix E, we elaborate on our implementation details for the encoder and decoder networks and the comparison settings to previous work.
@@ Appendix A Additional Results
Fidelity
iSeg is flexible and not bounded to parts defined in a dataset. Instead, it finds unique semantic regions of the considered shape. In Fig. 19, we compare our segmentation results to shape parts of the COSEG dataset. iSeg selects finer regions with high fidelity to the user’s input. For example, a click on the horse’s head segments only the head. The corresponding part in COSEG is coarser and includes both the head and neck. For the alien shape, all three antennas are regarded as one part. In contrast, iSeg marks only the antenna with the point click.
View generalization
Fig. 17 present additional examples of the view generalization of iSeg. Our method segments the whole body of the guitar, even though its back side is occluded from the click on the front. Moreover, the second click on the back side of the guitar’s bridge cannot be seen together in any 2D view with the first click. Nevertheless, our method selects the complete 3D regions of the bridge and body.
3D consistency
Different from SAM (Kirillov et al., 2023), our method operates natively in 3D. iSeg predicts the segmentation probability at the vertex level, or in other words, our result lives on the 3D surface of the mesh. Thus, it is 3D-consistent by construction. In Fig. 18, we compare our segmentation projected to 2D with SAM’s result for the projected shape views. Since SAM operates on each view independently, it yields inconsistent predictions. By contrast, in our case, any 2D projection originates from the same 3D segmentation and thus, our method is consistent across all views.
Separability
We analyze iSeg’s segmentation prediction by plotting the histogram of probability values for all the shape vertices for different clicks. Results were presented in Fig. 9 in the main paper. We observe that the histogram contains two populations: vertices with high segmentation probability, which correspond to the region for the clicked point, and the rest of the vertices with low probability. To further demonstrate the separability of the model’s predictions, we use the Otsu Threshold (Ostu, 1979) and binarize the probabilities. As the figure has shown, the region before and after thresholding remains almost the same, indicating that the segmented part can be easily separated from the shape.
Comparison
We present additional comparison results on the PartNet dataset in Fig. 23. While other methods struggle to segment the shape part corresponding to the click, our method succeeds.
@@
Full-shape segmentation
Our method’s purpose is to select a customized shape region corresponding to user clicks. Nonetheless, we can easily employ iSeg for parsing the entire shape into multiple parts. In this setting, the user applies several interactive segmentation sessions. Assuming we have $K$ sessions, in each one of them we get the vertex segmentation probability $p_{i}^{k}$, where $i$ is the vertex index and $k=1,...,K$ is the session index. Then, the vertex class $c_{i}$ is determined according to the maximal confidence:
@@
Figure 19. Segmentation fidelity. Our method selects fine-grained regions with higher fidelity to the clicked point compared to the segment defined in the COSEG dataset (van Kaick et al., 2011) that contains the clicked point.
(8) $c_{i}=\begin{cases}\operatorname*{\arg\!\max}\limits_{k}p_{i}^{k},&\text{if $\max\limits_{k}p_{i}^{k}\geq t_{seg}$}.\\ 0,&\text{otherwise}.\end{cases}$
$t_{seg}$ is a threshold for assigning a vertex to one of the segmentation regions, and class “0” is an additional background class. iSeg is typically confident in its predictions (with segmentation probabilities close to 00 or $1$), thus, we set $t_{seg}$ to $0.5$.
Fig. 10 presents examples of complete shape segmentation with several single-click sessions. iSeg’s region prediction for clicks on different parts of the share are localized and mostly disjoint from each other. It enables a clear separation of the entire object into its unique parts without the need for any part-label supervision.
Local geometric editing
In Fig. 13 in the main body, we show how our method’s results are used for local shape editing. iSeg selection corresponds to a contiguous region of an entire part of the shape. Thus, the segmented part can be easily manipulated. As the figure demonstrates, the entire bar of the chair is removed, only the head of the horse can be selected, the neck of the guitar is shortened, and a small local region of the candle’s flame is enlarged.
[…]
Figure 20. Segmentaton of close body parts. Although iSeg processes the mesh in the Euclidean space, it can separate body parts that are close in the Euclidean sense, but geodesic far.
Mesh processing in Euclidean space
Our model operates on the vertex coordinates directly in Euclidean space without considering the shape topology explicitly. Such processing of the mesh, or surfaces in general, may potentially result in features that do not disambiguate surface points that are close in Euclidean space but are distant geodesically. Nonetheless, in practice, iSeg manages to disambiguate Euclidean nearby shape segments.
As demonstrated in Fig. 20, when the human’s hand or arm is close to the torso, or the head is close to the hand, only the part corresponding to the clicked point is selected. These results suggest that our method has learned different semantic features for different shape parts from the 2D foundation model (Kirillov et al., 2023), enabling their separation. However, if the 2D supervision model does not disambiguate such regions successfully, iSeg may not distinguish them as well.
Stability
A desired property of a click-based segmentation model is to be stable. That is, the results should be similar for nearby clicks that relate to the same shape part. This property implies that the model is robust to the exact click location and alleviates the need for precision from the user. Fig. 21 demonstrates the stability of our method. As the figure shows, iSeg consistently selects the handle part of the hammer for different clicks along the handle.
We evaluate the stability property quantitatively by comparing the segmentation prediction for neighboring vertices. Given a clicked point, we compute the Intersection over Union (IoU) for its resulting mask and the mask for clicking each of its one-ring neighbors. We use 7 meshes with 100 random clicks from each mesh. For comparison, we utilize the SAM baseline, as described in Sec. 4.2 in the main paper. The average IoU for iSeg is 0.8 compared to 0.7 for the SAM baseline. SAM’s masks originate from 2D views of the shape. These predictions do not segment occluded regions of the shape and are less smooth in 3D. In contrast, our method operates directly in 3D and exhibits better stability than the SAM baseline.
In Fig. 24, we present the segmentation stability of iSeg. For a couple of clicks, the user can select the entire trailer box or only its side with various locations for the positive clicks or the positive and negative clicks, respectively.
Granularity
Fig. 22 exemplifies an additional property of our method: segmentation granularity. When segmenting the leg for an intricate shape like the human body, the user can control whether to exclude the foot, the lower shin part, or the upper part of the calf muscle by simply clicking the negative point at different locations along the leg.
Appendix B Quantitative Evaluation
We provide additional details about our quantitative evaluation on the PartNet dataset (Mo et al., 2019) discussed in Sec. 4.2 in the main paper. The PartNet dataset includes meshes with ground-truth part labels. We measured the segmentation performance of iSeg and other methods by clicking a vertex and comparing the results to the ground-truth annotation of the part containing the click. We did not use any ground-truth information for training, but rather only for performance evaluation.
For a robust evaluation, we randomly selected 5 vertices from the part that were not used during training. Each vertex was regarded as a single-click interactive segmentation session, and we averaged the results over the clicked points and the evaluation shapes. We employed meshes from the dataset’s test split for all the evaluated categories. According to our available resources, we used the 77 test shapes of the Knife category and sampled 93 additional meshes from the other categories.
B.1. Segmentaiton Binarization
iSeg predicts for each mesh vertex a soft segmentation value $p_{i}\in[0,1]$. However, the part labels in the PartNet dataset are categorial. Thus, we binarized our prediction to obtain a per-vertex segmentation result $q_{i}\in\{0,1\}$. The binarization was done with the Otsu Threshold (Ostu, 1979):
(9) $q_{i}=\mathbbm{1}(p_{i}\geq t_{otsu}),$
where $\mathbbm{1}{(\cdot)}$ is the indicator function and $t_{otsu}$ is the Otsu Threshold. We chose this simple binarization approach to minimize the post-processing effect and reflect our method’s performance as faithfully as possible. For a fair comparison, we applied the same binarization approach to the other baselines in our evaluation.
B.2. Evaluaiton Metrics
PartNet shapes are composed of mesh elements for each part with duplicate vertices at the elements’ connection. To circumvent this duplication, we converted the vertex prediction to face prediction $s_{i}$, where the face prediction was considered as ``$1"$ when all its vertices were segmented, and ``$0"$ otherwise. To compute the Accuracy (ACC) and Intersection over Union (IoU) metrics, we labeled the part that included the clicked point as ``$1"$, and the rest of the shape as label ``$0"$. The ACC measured the rate of prediction $s_{i}$ that matches the face segmentation label $l_{i}$:
(10) $ACC=\frac{1}{n}\sum_{i=1}^{n}{\mathbbm{1}(s_{i}=l_{i})}.$
For computing the IoU, we divided the intersection of the segmentation prediction and the ground-truth annotated region by their union:
(11) $IoU=\sum_{i=1}^{n}{(s_{i}\land l_{i})}\;\Big{/}\,\sum_{i=1}^{n}{(s_{i}\lor l_{i})}.$